# Learning to Generate Unit Tests for Automated Debugging

### v1

@@ 4 Experimental Setup
For additional details on these datasets including the number of test cases per problem and pass rate of faulty codes, we refer interested readers to Appendix A.
Evaluation Metrics.
@@
Code Accuracy. In order to evaluate the utility of generated unit tests via code debugging, we follow prior work (Chen et al., 2023b; Chae et al., 2024) in reporting pass@1 code accuracy, i.e., the percentage of codes passing all unit tests after 3 rounds of debugging. Note that while we debug with model-generated UTs, we evaluate code accuracy using private human-annotated UTs.
As the downstream application of unit-test generation is independent of the inherent task, we rely on intrinsic metrics during development on a validation split from MBPP+ data, e.g., for checkpoint selection and prompt design. Then we measure the utility and effectiveness of UTGen by comparing against baselines on code debugging tasks.
Baselines.
We compare UTDebug with feedback from UTGen against the following (prompts in Appendix E):
No UT feedback: In this baseline, we use self-generated or self-critique feedback, following prior work (Madaan et al., 2024; Shinn et al., 2024; Chen et al., 2023b). Specifically, we prompt the model to generate an explanation about the bugs in the target code in addition to determining if the target code is correct, i.e., does not have any more bugs (consistent with Chen et al. (2023b)).
Prompted Failing UTs: This baseline uses the same prompts as UTGen without any training and prompts an LLM to generate a failing unit test (with a correct output) given the problem description and the target code. Here, we sample .
5 Results and Analysis
c c c c c
Model Method Attack Rate Out Acc. Acc. Attack
Llama3 8B Random 30.98 46.86 10.24
Prompted 38.04 37.10 11.51
UTGen 41.18 48.24 16.57
Llama3.1 8B Random 21.76 39.80 9.02
Prompted 40.59 26.86 9.61
UTGen 41.37 47.75 16.67
Qwen2.5 7B Random 26.27 57.25 13.53
Prompted 39.80 40.78 14.71
UTGen 41.96 58.04 20.20
5.1 Intrinsic Evaluation of Unit Test Generation
In order to study the inherent UT generation abilities of different models and baselines, we use the intrinsic metrics defined by RQ1 and outlined in Section 4 in Section 5 on the most challenging debugging task MBPP+Fix (Hard) averaged over 3 runs.
Trade-offs in Attack Rate and Output Accuracy.
In Section 5, we benchmark the zero-shot abilities of different models on UT generation, as well as our improved training method (corresponding to RQ2). Here, we observe a clear tradeoff: prompted and random unit test generation each optimizes output accuracy or attack rate at the cost of the other metric. While randomly sampled UTs have relatively higher output accuracy (i.e., the output of the unit test is correct according to the problem description), the random baseline often lacks the ability to generate UTs with inputs that trigger errors (failing UTs). This can be explained by the fact that it is not conditioned on the faulty code . On the other hand, when models are prompted to generate a UT that breaks the faulty code in the prompted UT baseline, they generally achieve a higher attack rate; however, they lag in terms of output accuracy. For instance, in the case of Qwen2.5, switching from randomly sampled to prompting failing UTs improves the attack rate by 13.53% but decreases the output accuracy by 16.47%, hence leading to comparable “Acc. Attack” scores (13.53% vs 14.71%). Taken together, this suggests the presence of competing desirable traits of UT generators (cf. Section 3.1; RQ1): without training, one baseline often generates trivial unit tests for which the output can easily be predicted (high output accuracy, low attack rate), while the other generates challenging unit tests for which it cannot produce the expected answer (low output accuracy, high attack rate). It also indicates that while models can generate challenging unit test inputs that result in code failure, models struggle with reasoning over correct outputs of failing UT inputs.
UTGen is Most Effective at Unit Test Generation.
We propose to break through this trade-off (thereby addressing RQ2, “How can we improve LLMs’ UT generation abilities?”) by training models with UTGen, where we directly supervise models to have both high output accuracy and high attack rate. Ideally, UTGen should lead to models that generate unit tests in the “sweet spot” of difficulty, where they are hard enough to trigger errors but not so hard that their output cannot be predicted. Section 5 shows that models trained with UTGen consistently rank highest at jointly generating failing UT inputs with correct outputs (as measured via “Acc. Attack” in Section 5). For instance, on the strongest coding model in our setting, Qwen2.5, UTGen obtains the highest attack rate, and output accuracy as well as improves Acc. Attack score by 5.49% (absolute) over failing UTs generated by prompting. Similarly, Llama3.1 improves (in terms of attack rate) over randomly-sampled UTs by 19.61% and the output accuracy of prompted failing UTs by 20.89%, ultimately improving the joint “Acc. Attack” score by up to 7.65%. We report the intrinsic metrics for HE+Fix and MBPP+Fix in Appendix D.
5.2 Impact of Generated Unit Tests on Debugging
| Model | UT Method | HE+Fix | MBPP+Fix | (Hard) |
| Llama-3 8B | No UT | 27.22 | 16.31 | 11.76 |
| Random | 51.90 | 30.46 | 17.06 | |
| Prompted | 51.90 | 28.92 | 22.94 | |
| UTGen | 53.80 | 37.54 | 28.82 | |
| Llama-3.1 8B | No UT | 31.65 | 10.15 | 11.18 |
| Random | 62.03 | 33.54 | 13.53 | |
| Prompted | 56.33 | 28.00 | 24.71 | |
| UTGen | 67.09 | 36.92 | 28.23 | |
| Qwen-2.5 7B | No UT | 52.53 | 23.08 | 16.47 |
| Random | 79.75 | 34.77 | 17.06 | |
| Prompted | 75.32 | 32.92 | 24.12 | |
| UTGen | 82.91 | 37.54 | 29.41 |
UTs from UTGen are Best for UTDebug.
Addressing RQ3, we measure how effective each type of UT generator is in a downstream debugging evaluation. From Table 2, we observe that across different LLM families, multi-turn debugging with UTDebug is more effective when using generated UTs from UTGen than debugging with UTs generated by the baselines, as well as debugging without feedback. For instance, with Qwen2.5, after 3 rounds of debugging on MBPP+Fix, UTGen improves over the randomly-sampled UT baseline by 3% (absolute), debugging by failing UTs by 7.59% and without UT feedback by 21.23% (absolute). Moreover, on the more challenging MBPP+Fix (Hard) split, UTGen improves over randomly-sampled UTs by 12.35% and the baseline without UT feedback by 12.94%. Given that we use the same underlying LLM and similar feedback templates, the results in Table 2 show that UTs generated by UTGen provide the most useful and effective feedback. Lastly, we find that LLM debugging without any UT feedback is least effective across model families, thus, establishing that despite noise and issues discussed in Section 3.3 and Figure 3, UTDebug with UT-based feedback is more effective than self-generated feedback. This aligns with past work indicating that LLMs are poor at self-critique (Huang et al., 2024; Stechly et al., 2024; Chen et al., 2024).
Debugging Difficulty Varies Across Datasets. Comparing the post-debugging accuracy across the three datasets in Table 2 suggests different difficulty levels when it comes to LLM debugging. The relatively high accuracies on HE+Fix (especially with Qwen-2.5) suggest that human-introduced errors in that dataset are obvious and thus too easy for the models to discover and debug. This is also corroborated by the fact that starting codes in this dataset have a pass rate of 0%, i.e., fail on all unit tests, suggesting high-level coarse-grained errors, e.g., syntax errors that prevent execution, or incorrect function references. On the other hand, MBPP+Fix (Hard) (with an initial pass rate of 75%) appears to be the hardest to debug, with the lowest overall post-debugging accuracy across models – up to a difference of 53.5% in final accuracy for UTGen based on Qwen-2.5. This in turn suggests that LLMs still struggle with identifying and fixing subtle flaws in generated code, especially in scenarios involving corner cases wherein the model passes some unit tests and fails at the rest.
5.3 Effectiveness of the UTDebug Pipeline
In Section 3.3, we highlighted the challenges of debugging with imperfect model-generated unit tests and suggested remedies to make our debugging pipeline, UTDebug, robust to noisy feedback from such UTs. We study the effectiveness of these measures, i.e., test-time scaling of UT outputs, and backtracking, on debugging with generated UTs for 3 rounds using Qwen2.5 on the MBPP+Fix dataset in Table 3.
| UT Method | Acc. | |
| Randomly-sampled (Qwen2.5) | 34.77 | |
| - Output Test-time Scaling | 30.77 | - 4.0 |
| - Backtracking | 32.61 | - 2.2 |
| UTGen (Qwen2.5) | 37.54 | |
| - Output Test-time Scaling | 26.15 | - 11.4 |
| - Backtracking | 34.38 | - 3.2 |
Test-time Scaling and Backtracking are Crucial for LLM Debugging with Generated UTs. From Table 3, we observe that irrespective of the underlying method for UT generation (either randomly sampled or from UTGen), removing either backtracking or test-time scaling hurts downstream performance. First, we find that removing test-time scaling for predicting the output of UTs decreases the performance of randomly-sampled UTs by 4% and that of UTGen by 11.4%. Note that a larger drop in UTGen’s performance when removing test-time scaling of UT output prediction is consistent with the findings in Section 5.1 that models struggle with reasoning over correct outputs for failing UT inputs (more often generated by UTGen). Therefore, test-time scaling in output prediction provides a greater boost for reasoning over challenging inputs (Wang et al., 2022), and consequently, removing it yields larger drops for UTGen. Additionally, Table 3 demonstrates that, without backtracking and validation on other generated unit tests, LLMs tend to overfit on the unit test contained in the feedback, resulting in up to 3.2% drop in performance of UTGen without backtracking.
5.4 Scaling with Number of Generated UTs
Thus far, all our experiments use generated UTs across baselines and models. However, as we described in Section 3.3, having multiple UTs can be advantageous because: (i) there is a higher likelihood of generating a failing UT and getting a reliable signal for when the code is correct, (ii) more robust signal for when to backtrack using validation on the entire generate test suite. We analyze the impact of increasing the number of generated unit tests on downstream accuracy of Qwen2.5 after 3 rounds of debugging in Figure 4. Our findings are as follows:
First, Figure 4 shows that despite increasing the number of generated unit tests, UTGen consistently outperforms randomly sampled UTs that may not be failing. This highlights the benefits of generating targeted unit tests conditioned on buggy code in order to trigger errors and generate appropriate feedback for debugging.
In the case of MBPP+Fix (Hard), which contains codes with less obvious errors and are harder to debug, we find that despite scaling to up to generated UTs, the performance gap between UTGen and randomly-sampled UTs remains at 10% (absolute).
5.5 Comparison to Frontier LLMs
c c c c c
Model Method Attack Rate Out Acc. Acc. Attack
Qwen2.5 7B Random 26.27 57.25 13.53
Prompted 39.80 40.78 14.71
UTGen 41.96 58.04 20.20
GPT-4o Random 23.92 57.25 17.25
Prompted 54.51 60.39 33.33
DeepSeek3 Random 24.31 63.53 17.84
Prompted 54.31 58.04 33.33
In Section 5 we evaluated open-source LLMs that we trained using UTGen, and in Section 3.3 we applied these models in our UTDebug framework for debugging, finding the trained models to be effective. Section 5.5 compares stronger frontier models – GPT-4o (Achiam et al., 2023) and DeepSeek-V3 (Guo et al., 2024b) – to the best open-source 7-8B model tested (Qwen2.5 7B). Here, we see that, while they outperform Qwen, even these stronger and much larger frontier models struggle at the task of generating failing unit tests. First, we find that when prompting models with faulty codes, the attack rate of the generated UTs is slightly over 50% on MBPP+Fix (Hard) dataset, showcasing the inherently challenging nature of isolating corner cases and subtle errors in partially correct code. Furthermore, both GPT-4o, DeepSeek-V3 attain comparable output accuracies to that of our trained UTGen model on top of Qwen2.5, while being larger and more costly to run (although their Acc. Attack scores are higher). Overall, we find that using UT-based feedback even with stronger LLMs only triggers and identifies errors 1 in 3 times, leaving tremendous room for growth in the ability of models to identify and provide feedback on partially correct code.
Note that cost is a major factor in automated debugging; we show that using UTGen for debugging in UTDebug is successful, but involves generating multiple UTs with self-consistency across several rounds of debugging for each problem (cf. Algorithms 1 and 2 in Appendix C). Such calls will quickly become costly on frontier models like GPT-4o and DeepSeek-3, making 7-8B models like the ones we train with UTGen attractive options. Moreover, since these frontier models are not trainable, there is less room for improvement when scaling training data.
6 Discussion and Conclusion
Our work on UTGen contributes to the broader landscape of verification and feedback generation for LLM-generated code. While recent work has focused on training verifiers to provide feedback (Mahan et al., 2024; Zhang et al., 2024), a key challenge remains in obtaining high-quality feedback signals for debugging. UTGen addresses this by directly generating unit tests that can identify problems in code, complementing existing work on how to effectively incorporate and present feedback for debugging (Chen et al., 2023b; Zhong et al., 2024) along with test-time scaling and backtracking incorporated in UTDebug. Our results demonstrate that without a quality signal to determine code correctness and/or how a faulty code is failing (in the form of unit tests), using LLMs to generate feedback and debug still proves to be challenging. This is one of the first efforts in this direction, and we hope to spark more interest in future work toward LLM-generated unit tests (both input and outputs) that reveal the model’s coding errors.
Our approach connects to and complements recent work on handling real-world programming issues. While approaches designed for SWEBench (Jimenez et al., 2024) focus on fixing known issues from GitHub by understanding and implementing fixes for bug reports, UTGen addresses the upstream challenge of automatically discovering potential issues in new code through test generation. Both tasks share a core challenge: determining the expected behavior of code without access to correct implementations. This connects to the fundamental concept of simulatability from computability theory (Rice, 1953), where we ask whether a program can predict the behavior of another program. Recent work such as Jain et al. (2024) shows that while LLMs can often simulate existing code by tracing execution steps, they struggle more with predicting correct outputs from specifications alone. Our results align with these findings – while UTGen can generate test inputs that trigger errors (high attack rate), predicting correct expected outputs remains challenging (lower output accuracy). This suggests that improving LLMs’ ability to reason about intended program behavior from specifications remains a crucial direction for future work. Nevertheless, we find that the modifications made to debugging in UTDebug help boost UTGen’s accuracy and account for noise, leading to downstream gains.
Conclusion.
We first identified a key trade-off between attack rate and output prediction accuracy when predicting unit tests with models. In light of this trade-off, we introduced UTGen, a new method for creating training data and teaching models to produce unit tests. This allows us to train models to produce better unit tests, as measured by intrinsic metrics like attack rate and output accuracy. Moreover, finding that existing data contains large numbers of easy errors, we introduce a new subset of data with challenging and hard-to-diagnose errors. To enable debugging with automated unit tests, we propose UTDebug, wherein we augment our predictor’s accuracy with test-time scaling and regularize it using a cross-validation and back-tracking procedure that prevents it from overfitting to a narrow or incorrect unit test. This, combined with our improved predictor, results in consistent increases in debugging performance across models, a gap that persists as we scale the number of unit tests.
Acknowledgments
We thank Swarnadeep Saha for insightful discussions in the early phases of this work. This work was supported by NSF-CAREER Award 1846185, DARPA ECOLE Program No. HR00112390060, and NSF-AI Engage Institute DRL-2112635. Any opinions, findings, and conclusions or recommendations in this work are those of the author(s) and do not necessarily reflect the views of the sponsors.
Appendix A Debugging Datasets
HE+Fix. This dataset contains a total of 158 problems each with one incorrect human-written code and has an initial pass rate (prior to debugging) of 0%, i.e., all private unit tests are failing. As mentioned in Section 4, we use the dataset provided by Muennighoff et al. (2024)444https://huggingface.co/datasets/bigcode/humanevalpack but replace the test set for each problem from the original test suite in HumanEval (Chen et al., 2021) to that in the EvalPlus evaluation suite (Liu et al., 2024). This increases the average unit tests per problem from 8.17 to 775.18 gold unit tests. Note that we have an automatic UT extraction script for the test code in EvalPlus, and we only retain problems for which this extraction is successful (158 out of 164).
MBPP+Fix and MBPP+Fix (Hard). We begin with 378 problems in the MBPP+ dataset (Liu et al., 2024)555https://huggingface.co/datasets/evalplus/mbppplus and follow the same gold UT extraction step described above, discarding problems for which the extraction fails. This leaves us with 375 problems, for which we sample 16 solutions per problem across multiple LLMs: Gemma-7B-IT (Team et al., 2024), Llama3 8B Instruct (AI@Meta, 2024), Llama3.1 8B Instruct (Dubey et al., 2024), Tulu-3 8B SFT (Lambert et al., 2024), DeepSeek 7B coder (Guo et al., 2024b), Qwen-2.5 Coder (Hui et al., 2024). To generate MBPP+Fix, we filter for incorrect solutions (i.e., with at least one failing UT) and then randomly sample one incorrect code per problem. This yields 325 problems in total, each with one faulty code. This dataset has an initial pass rate of 24.21% and an average of 107.45 gold unit tests per problem. In order to construct, MBPP+Fix (Hard), we follow a similar process but select only incorrect solutions which pass 50 - 95% of unit tests. The intuition here is that solutions that are partially correct are often harder to debug than those that are fully incorrect. We then randomly sample one such incorrect solution per problem, yielding a dataset of 170 problems with an initial pass rate of 74.83% and an average of 107.49 gold unit tests.
@@ Appendix B UTGen Training
Preprocessing Tulu Data. We use the Tulu-3 SFT mixture dataset released by Lambert et al. (2024) which contains a total of 939.3K instances.666https://huggingface.co/datasets/allenai/tulu-3-sft-mixture However, it contains a mixture of prompts for instruction-following, math reasoning, and coding. Therefore, we filter for prompts involving Python coding by regex search for keywords “python” and “def ” which suit our task setup described in Section 3.1. Furthermore, we filter out instances with more than 2K tokens in the prompt and ensure the prompt is a valid unicode string. This results in a total of 48.3K instances for which we use the “prompt” as the problem description and extract code from the response of the last turn when multi-turn interactions are present provided that the extracted code executes without any errors or issues. Finally, we prompt the LLM to be trained with UTGen to generate 2 corrupted versions of this code to serve as the target code, and for each target code, we make 5 attempts to generate failing unit test inputs and filter out instances that do not have any such UTs. This is followed by the rationalization step (with the same underlying LLM to be trained) described in Section 3.2 and in Figure 2, which results in roughly 30K instances for each model trained with UTGen (Qwen2.5, Llama-3, -3.1 8B).
Training Hyperparameters. We train each model for 10 epochs with a batch size of 16 and a learning rate of 5e-6 with a cosine learning rate scheduler. Moreover, we only compute negative log-likelihood loss over the completions. We use LoRA (Hu et al., 2021) with a rank of 16, = 32 with a dropout of 0.05. We perform checkpoints selection based on the intrinsic Acc. Attack metric.
@@ Appendix C Overall Pipeline for UTDebug
In Algorithm 1 we describe the process of generating UTs for a candidate buggy code using any UT generation method and perform test time scaling for UT output prediction. This is then used within UTDebug as shown in Algorithm 2, which identifies failing UTs, debugs the target code based on feedback from failing UTs over multiple rounds, and returns the debugged (edited code). As illustrated in Figure 3 (b), edits are accepted only if the newly generated code achieves a higher pass rate than the code before debugging, otherwise the edits are discarded.
@@ Appendix D Additional Intrinsic Evaluation
Similar to Section 5, we report intrinsic evaluation metrics (cf. Section 4) for HE+Fix and MBPP+Fix datasets in Table 5. Consistent with the findings in Section 5.1, without training, we observe a trade-off between attack rate and output accuracy, with randomly sampled UTs showing higher output accuracy whereas the prompted baseline exhibits higher attack rates and vice-versa. Once again UTGen training best balances this trade-off yielding both high output accuracy and attack rate. However, due to the relative ease of attacking faulty codes in HE+Fix (initial pass rate of 0%) almost any generated UT fails and thus can be used for debugging. This, coupled with the higher output accuracy, results in the random baseline having the highest score on Acc. Attack. Note that we mitigate this impact on downstream performance by using test-time scaling for output prediction in UTDebug, which especially boosts the accuracies of targeted UTs generated based on the buggy code. On MBPP+Fix, UTGen consistently yields the highest Acc. Attack score, followed by the random baseline.

### v2

@@ A.1 Coding Datasets
HE+Fix. This dataset contains a total of 158 problems each with one incorrect human-written code and has an initial pass rate (prior to debugging) of 0%, i.e., all private unit tests are failing. As mentioned in Sec. 4, we use the dataset provided by Muennighoff et al. (2024)444https://huggingface.co/datasets/bigcode/humanevalpack but replace the test set for each problem from the original test suite in HumanEval Chen et al. (2021) to that in the EvalPlus evaluation suite Liu et al. (2024b). This increases the average unit tests per problem from 8.17 to 775.18 gold unit tests. Note that we have an automatic UT extraction script for the test code in EvalPlus, and we only retain problems for which this extraction is successful (158 out of 164).
MBPP+Fix and MBPP+Fix (Hard). We begin with 378 problems in the MBPP+ dataset Liu et al. (2024b)555https://huggingface.co/datasets/evalplus/mbppplus and follow the same gold UT extraction step described above, discarding problems for which the extraction fails. This leaves us with 375 problems, for which we sample 16 solutions per problem across multiple LLMs: Gemma-7B-IT Team et al. (2024), Llama3 8B Instruct AI@Meta (2024), Llama3.1 8B Instruct Dubey et al. (2024), Tulu-3 8B SFT Lambert et al. (2024a), DeepSeek 7B coder Guo et al. (2024b), Qwen2.5 Coder 7B Hui et al. (2024). To generate MBPP+Fix, we filter for incorrect solutions (i.e., with at least one failing UT) and then randomly sample one incorrect code per problem. This yields 325 problems in total, each with one faulty code. This dataset has an initial pass rate of 24.21% and an average of 107.45 gold unit tests per problem. In order to construct, MBPP+Fix (Hard), we follow a similar process but select only incorrect solutions which pass 50 - 95% of unit tests. The intuition here is that solutions that are partially correct are often harder to debug than those that are fully incorrect. We then randomly sample one such incorrect solution per problem, yielding a dataset of 170 problems with an initial pass rate of 74.83% and an average of 107.49 gold unit tests.
Code Generation.
When using generated UTs for improving code generation, we evaluate on HumanEval+ (HE+) and MBPP+ benchmarks proposed in Liu et al. (2024b) released under Apache-2.0 license. As described above, we filter out problems for which we can extract private UTs, yielding 158 problems in HE+ dataset and 375 problems in MBPP+ dataset. For the RM baseline we use the latest version of the 8B model based on Llama-3.1 backbone in Liu et al. (2024a).666https://huggingface.co/Skywork/Skywork-Reward-Llama-3.1-8B-v0.2 We made this choice based on on rankings on the RewardBench leaderboard Lambert et al. (2024b)
A.2 Evaluation Metrics
@@
Code Accuracy. To evaluate the utility of generated unit tests via code debugging, we follow prior work Chen et al. (2023b); Chae et al. (2024) in reporting pass@1 code accuracy, i.e., the percentage of codes passing all unit tests after 3 rounds of debugging. Note that while we debug with model-generated UTs, we evaluate code accuracy using private human-annotated UTs.
@@ Appendix B UTGen Training
Preprocessing Tulu Data. We use the Tulu-3 SFT mixture dataset released by Lambert et al. (2024a) which contains a total of 939.3K instances.777https://huggingface.co/datasets/allenai/tulu-3-sft-mixture However, it contains a mixture of prompts for instruction-following, math reasoning, and coding. Therefore, we filter for prompts involving Python coding by regex search for keywords “python” and “def ” which suit our task setup described in Sec. 3.1. Furthermore, we filter out instances with more than 2K tokens in the prompt and ensure the prompt is a valid unicode string. This results in a total of 48.3K instances for which we use the “prompt” as the problem description and extract code from the response of the last turn when multi-turn interactions are present provided that the extracted code executes without any errors or issues. Finally, we prompt the LLM to be trained with UTGen to generate 2 corrupted versions of this code to serve as the target code, and for each target code, we make 5 attempts to generate failing unit test inputs and filter out instances that do not have any such UTs. This is followed by the rationalization step (with the same LLM to be trained) described in Sec. 3.2 and in Fig. 2, which results in roughly 30K instances for each 7-8B model trained with UTGen and nearly 70K instances for Qwen2.5 32B.
@@
| Qwen2.5 32B | Random | 84.96 | 80.19 | 79.95 | 54.56 | 50.17 | 34.87 |
| Prompted | 88.82 | 64.77 | 63.89 | 66.87 | 35.59 | 27.38 | |
| UTGen | 95.93 | 70.64 | 69.79 | 70.36 | 50.87 | 40.62 | |
Training Hyperparameters.
We train each model for 10 epochs with a batch size of 16 and a learning rate of 5e-6 with a cosine learning rate scheduler. Moreover, we only compute negative log-likelihood loss over the completions. We use LoRA Hu et al. (2021) with a rank of 16, = 32 with a dropout of 0.05. We perform checkpoint selection based on the intrinsic Acc. Attack metric. All training and inference is conducted on Nvidia’s A6000 GPUs taking 20 GPU hours for training and GPU hours for inference (generating one code solution and corresponding UTs with test-time scaling). We train the larger Qwen2.5 32B model on 8 A100 GPUs for roughly 40 GPU hours using QLoRa Dettmers et al. (2023) with 4-bit quantization.
@@ Appendix C Overall Pipeline for UTDebug
In Algorithm 1 we describe the process of generating UTs for a candidate buggy code using any UT generation method and perform test time scaling for UT output prediction. This is then used within UTDebug as shown in Algorithm 2, which identifies failing UTs, debugs the target code based on feedback from failing UTs over multiple rounds, and returns the debugged (edited code). As illustrated in Fig. 3 (b), edits are accepted only if the newly generated code achieves a higher pass rate than the code before debugging, otherwise the edits are discarded.
@@ Appendix D Additional Intrinsic Evaluation
Similar to Sec. 5.1, we report intrinsic evaluation metrics (cf. Sec. 4) for HE+Fix and MBPP+Fix datasets in Table 6. Consistent with the findings in Sec. 5.1, without training, we observe a trade-off between attack rate and output accuracy, with randomly sampled UTs showing higher output accuracy whereas the prompted baseline exhibits higher attack rates and vice-versa. Once again UTGen training best balances this trade-off yielding both high output accuracy and attack rate. However, due to the relative ease of attacking faulty codes in HE+Fix (initial pass rate of 0%) almost any generated UT fails and thus can be used for debugging. This, coupled with the higher output accuracy, results in the random baseline having the highest score on Acc. Attack. Note that we mitigate this impact on downstream performance by using test-time scaling for output prediction in UTDebug, which especially boosts the accuracies of targeted UTs generated based on the buggy code. On MBPP+Fix, UTGen consistently yields the highest Acc. Attack score, followed by the random baseline.